OpenStudy (anonymous):

In a 45°-45°-90° triangle, the length of the hypotenuse is 11. Find the length of one of the legs.

OpenStudy (johnweldon1993):

A 45 45 90 triangle has the other 2 legs of the same length... and the hypotenuse...is the length of 1 of those legs...times √2 So we need to solve \[\large \sqrt{2}x = 11\]

OpenStudy (anonymous):

1.41x=11?

OpenStudy (johnweldon1993):

Well I wouldnt approximate just yet...I would simply do \[\large x = \frac{11}{\sqrt{2}}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@johnweldon1993 do you know how i would do this problem? Which of the following are not the lengths of the sides of a 30°-60°-90° triangle?

OpenStudy (johnweldon1993):

Sure... So for that one...the side ratios are 1:2:√3  So whatever the shortest side is...make that X Then the hypotenuse will be 2 times that... And the length of the 3rd leg will be the smallest leg..times \(\sqrt{3}\)

OpenStudy (anonymous):

how do i start figureing out what x is

OpenStudy (johnweldon1993):

....judging from the question...you should have choices...

OpenStudy (anonymous):

\[A.\frac{ 1 }{ 2}, \frac{ \sqrt{3} }{ 2 },1 \]

OpenStudy (anonymous):

\[B.\frac{ 5 }{ 2 },\frac{ \sqrt[5]{?3} }{ 2},10\]

OpenStudy (anonymous):

\[C.\sqrt{2},\sqrt{6},\sqrt[2]{2}\]

OpenStudy (anonymous):

\[D.3,\sqrt[3]{3},6\]